# MapEx: Indoor Structure Exploration with Probabilistic Information Gain from Global Map Predictions

### v2

@@ Abstract
Exploration is a critical challenge in robotics, centered on understanding unknown environments. In this work, we focus on robots exploring structured indoor environments which are often predictable and composed of repeating patterns. Most existing approaches, such as conventional frontier approaches, have difficulty leveraging the predictability and explore with simple heuristics such as ‘closest first’. Recent works use deep learning techniques to predict unknown regions of the map, using these predictions for information gain calculation. However, these approaches are often sensitive to the predicted map quality or do not reason over sensor coverage. To overcome these issues, our key insight is to jointly reason over what the robot can observe and its uncertainty to calculate probabilistic information gain. We introduce MapEx, a new exploration framework that uses predicted maps to form probabilistic sensor model for information gain estimation. MapEx generates multiple predicted maps based on observed information, and takes into consideration both the computed variances of predicted maps and estimated visible area to estimate the information gain of a given viewpoint. Experiments on the real-world KTH dataset showed on average 12.4% improvement than representative map-prediction based exploration and 25.4% improvement than nearest frontier approach.
@@ I Introduction
Exploration is an important problem in robotics, with applications ranging from indoor search and rescue operations [1] to planetary missions [2] and outdoor monitoring [3]. The primary objective of exploration is to plan paths in an unknown environment to gain as much understanding of the environment within a limited budget. This objective could include building an accurate map [4], finding objects of interest [5], and finding traversable routes for robot teams [6]. The key challenge in exploration is choosing how to navigate for most information gain which requires reasoning over unobserved space due to uncertain environment geometry, occlusions, sensor range limits, and sensor noise.
Figure 1: We present MapEx, a new framework that uses predicted maps for indoor exploration. Our key insight is to calculate probabilistic information gain that jointly reasons on coverage and uncertainty. We validate with a real large academic building dataset and show superior exploration performance compared to related map-prediction based exploration algorithms.
Many real-world environments that we aim to explore have some kind of structure, predictability, and repeatability of environment geometry. For example, indoor environments like offices or hospitals consist of multiple repeated rooms and corridors; observing partial information of the space can help predict the overall layout of the rest of the environment. However, most existing approaches have difficulty leveraging the predictability in such environments for efficient exploration. For instance, methods like conventional frontier exploration [7, 8] do not directly reason over the predictable structure of the environment, and instead rely on heuristics such as “visit closest area first”.
To leverage such structural predictability, there has been recent work that uses deep learning techniques to predict unknown regions of maps, and use such prediction to improve exploration. For instance, IG-Hector [9] proposed using a predicted map to denote how much area can be observed from a viewpoint. However, such viewpoint scoring heuristic is sensitive to the predicted map accuracy. UPEN [10] instead leverages such uncertain map prediction behavior, by predicting map uncertainty explicitly and then going towards these areas. However, this method does not consider sensor coverage or visibility, which do not reflect actual sensor acquisition. Therefore, our key insight is to jointly reason over what the robot can observe and its uncertainty to form probabilistic information gain from multiple predicted maps.
@@
  • Finally, we show the predicted map produced by MapEx have improved global topological understanding and higher utility for downstream path planning.
@@ II-C Map Prediction for Exploration
To bring more expressiveness, deep learning emerges as a good approach for map prediction. Several works leverage deep learning-based map predictors with deep reinforcement learning-based planners for exploration [29, 30, 31]. However, due to the need for large amounts of data, these approaches are limited to short horizon decisions, such as 2.5m away from robot [29], one step towards a direction [30, 31].
Our work builds on successful approaches that leverage deep-learning map prediction techniques and couple them with classical exploration planning techniques that allows more explicit planning over longer horizons, and no need for data to learn planning [9, 37, 28, 32, 10]. One method to leverage the prediction for exploration is to estimate the potential new information from a predicted sensor coverage at a viewpoint and use it as an information gain metric [9, 37, 28, 32]. However, this information gain metric is susceptible to poor map predictions. Alternatively, UPEN [10] leverages such uncertainty in map prediction, to bias exploration towards high-uncertainty areas. UPEN estimates map uncertainty by predicting an ensemble of maps and using its variance as uncertainty. However, the information metric used is purely variance over the path and does not consider the potential sensor coverage, which is important to encapsulate actual information gain. To address the limitations of past work, MapEx jointly reasons over the predicted sensor coverage and the uncertainty of map prediction to form probabilistic information gain that is easily used by classical exploration frameworks [7], to generate long-horizon and informative exploration paths.
Figure 2: MapEx Framework: Our exploration approach estimates probabilistic information gain from predicted maps to use in a frontier explorer. First, a robot accumulates its observations into a top-down observed occupancy map which is then passed to an ensemble of map predictors to generate multiple map predictions. For each possible frontier, information gain is calculated from the mean and variance of the predicted maps. Finally, the frontier with the highest reward is chosen as the next goal.
@@ III Problem Statement
We address the problem of robot exploration in unknown indoor environments to gain a comprehensive understanding of the structure. Let’s assume that a robot is exploring in a 2D environment $\mathcal{E}\subset\mathcal{R}^{2}$. The robot’s state can be defined as a 2D discrete state $\mathbf{x}=[x,y]$, and the robot’s action $a\in\mathcal{A}$ transitions its pose in the navigation graph on a diagonally-connected grid, per each timestep. The robot is equipped with a 2D LiDAR sensor with range $\lambda$. At time $t=0$, the robot starts with no prior knowledge of the environment, and at each timestep $t$, it receives 2D LiDAR measurements $o_{t}$ from evenly spaced rays spanning $360^{\circ}$ originating from the robot state $\mathbf{x_{t}}$, and takes an action $a_{t}$ according to the planner. We assume that the robot’s pose in a global coordinate frame is known and noise-free LiDAR range measurements are available. It continues to take actions and receive measurements until the time budget $T$.
@@ IV Approach
To efficiently explore indoor structured environments, we leverage predicted maps to find high probabilistic information gain areas. The MapEx pipeline is shown in Fig. 2, consisting of four key steps:
@@ IV-A Global Map Prediction
Given the prevalence of large unknown areas during indoor exploration, we use LaMa network architecture [26], which has image-wide receptive field and ability to understand global context, making it suitable for inpainting maps with large unknown areas. We treat a top-down view occupancy grid as a 2D image (where 1 pixel = $0.1$m) and regard the unknown areas of the environment similarly to missing parts of an image. LaMa is used to predict what lies beyond the robot’s observation. To increase map prediction performance, we fine-tuned LaMa’s out-of-the-box Places2 weights using the KTH floorplan dataset[11]. Our training set consists of (observed map, ground truth map) data pairs collected through nearest-frontier exploration. Unlike prior works that predicts local maps [9, 10], we generate global predicted maps to enable more comprehensive global planning in exploration. Fig. 3 shows an example of an observed map, associated predicted map and ground-truth map.
@@ IV-B Uncertainty Estimation and Variance Map
While the robot can reasonably predict the map $P_{t}$ given the observed map $O_{t}$, there is inherent ambiguity in the unobserved areas. Therefore, the robot should estimate the uncertainty of the predictions. To achieve this, we maintain an ensemble of separate LaMa networks $\mathcal{G}_{i}$, and let each $\mathcal{G}_{i}$ generate prediction:
@@ IV-C Calculating Probabilistic Information Gain
Given a viewpoint $c\in\mathcal{E}$, we generate $n_{l}$ hypothetical rays $R_{j}$ (where $j=1,2..,n_{l}$) with a range of $\lambda$, emitted in a $360^{\circ}$ pattern originating from $c$. In a deterministic raycast approach, each ray $R_{j}$ stops when it hits an occupied cell, and if there are no obstructing cells along its path, $R_{j}$ reach the end at a distance of $\lambda$. In this work, we use a probabilistic raycast approach instead: first, we calculate the mean predicted map $\overline{P_{t}}$, which is the average of predicted maps $\{P_{i,t}\}$. Each ray $R_{j}$ starts with an initial accumulated occupancy value $\Delta=0$, and as $R_{j}$ passes through the pixels of $\overline{P_{t}}$, it adds the occupancy value of each pixel in the mean map $\overline{P_{t}}$ to $\Delta$. When $\Delta$ reaches a threshold $\epsilon$, the raycast terminates. Given ray end points, we perform a flood fill to generate a sensor coverage mask. We then mask out already observed areas from the sensor coverage mask to generate a visibility mask $\nu$, which probabilistically represents the unknown area estimated to be visible at location $c$. We then compute information gain I by summing the values of all pixels in the variance map $V_{t}$ that also fall within the visibility mask $\nu$. In other words,
@@ IV-D MapEx: Map Exploration with Prediction
Here, we explain how the MapEx algorithm is structured and uses the information gain values to augment a frontier-based explorer for efficient exploration. When the robot needs to select the next frontier, it generates a predicted map $P_{t}$ based on the current observed occupancy grid $O_{t}$, and a variance map $V_{t}$ and ensembles of predictions $\{P_{i,t}\}$. The frontiers are extracted from the boundary between known and unknown cells in $O_{t}$, similar in [7]. When calculating the score for these frontiers, the information gain is measured as described in Sec. IV-C. To discourage backtracking, we divide each frontier score by the Euclidean distance from the current pose of the robot to each frontier, so that if the scores are similar, the robot prefers to select a closer frontier. Once the waypoint is chosen, a local path is generated using an A* local planner, and the robot moves along the path. The pseudocode of the algorithm can be found in Alg. 1.
@@ V Experiments
We evaluate MapEx algorithm using a simulator on real-world floorplan datasets by comparing it with state-of-the-art baselines, and present ablation studies of the algorithm.
@@ V-A Real-world Floorplan Dataset
To benchmark real-world performance, we use the KTH floorplan dataset [11], which after cleaning repeated maps, contains 149 campus floorplan blueprints described as XML files of wall and door locations for each room. We adapt the processing code from [38] to obtain the occupancy map. In addition, we correct incorrect areas of the processed map. We then downsample the map to $0.1$m per pixel. For training the map predictor and then testing the navigation, we split the floorplans into an 80:20 train-test split. We take care to not include floorplans from the same building into both the train and test set, to test the generalizability of our method.
@@ V-C Baselines
We also ablate the impact of different components of the information gain metric used by MapEx, to quantify the relative impact. For all methods, we kept the frontier exploration framework the same, with only the frontier scoring method different. Specifically, we investigate
@@
  • No Visibility - No raycast is done. We sum the variance within $5$m of the viewpoint.
@@ V-D Metrics
We evaluate our MapEx planner using a variety of metrics to fully evaluate the information gained by the robot: coverage, predicted map quality, and topological understanding.
@@
Predicted Map Quality: The intersection over union (IoU) of the predicted 2D occupancy map and the ground truth occupancy map. Specifically, we use the IoU of the occupied class within the building footprint. Higher IoU means the predicted map is more accurate.
Topological Understanding: We propose a new metric, which we call Topological Understanding (TU), to directly assess how much the outcome of a robot’s exploration aids downstream path planning tasks. The TU metric is defined as follows: based on the exploration result up to timestep $t$, we have a predicted map $P_{t}$. Now we can plan an A* path from a starting point to an end-point goal. If the path successfully reaches the goal, it is considered a success. If the path collides with an occupied cell in the ground-truth map or fails to reach the goal, it is considered a failure. We measure success rates by sampling 100 random end-point goals for each map and start point across different methods.
@@ V-E1 Qualitative Analysis
Fig. 5 compares exploration progress over time for one test map using different methods. UPEN [10] reasons on variance at the path and does not consider sensor coverage, likely leading it to explore areas with high variance near the center of the map. In contrast, at $t=500$, MapEx also explores the center, but quickly shifts to a more informative hallway, estimating high sensor coverage. IG-Hector [9], which focuses on sensor coverage, struggles early likely due to poor map predictions, causing inefficient exploration of tight spaces. In comparison, as seen in $t=200$MapEx searches a relatively open area, gathering information about surrounding rooms. Overall, MapEx covers more informative areas within the same time frame, resulting in a better predicted map compared to the baselines. Fig. 6 provides more qualitative examples of the MapEx’s frontier scoring, showcasing the utility of jointly reasoning on sensor coverage and uncertainty. Fig. 6a,b show top-scoring frontiers with high sensor coverage with medium uncertainty. In contrast, Fig. 6c presents a scenario where MapEx selects a frontier that covers smaller area but has high uncertainty, rather than one with higher coverage but less potential for learning, highlighting the importance of reasoning about uncertainty.
@@ VII Acknowledgement
We thank Siddarth Narasimhan for sharing a more processed KTH dataset and providing valuable insights into baseline reproduction. We thank Simon Stepputtis, Yaqi Xie, David Fan and Mononito Goswami for helpful discussions.

### v3

@@ Abstract
Exploration is a critical challenge in robotics, centered on understanding unknown environments. In this work, we focus on structured indoor environments, which often exhibit predictable, repeating patterns. Conventional frontier-based exploration approaches have difficulty leveraging this predictability, relying on simple heuristics such as ‘closest first’ for exploration. More recent deep learning-based methods predict unknown regions of the map for information gain computation, but these approaches are often sensitive to the predicted map quality or fail to account for sensor coverage. To overcome these issues, our key insight is to jointly reason over what the robot can observe and its uncertainty to calculate probabilistic information gain. We introduce MapEx, a new exploration framework that uses predicted maps to form probabilistic sensor model for information gain estimation. MapEx generates multiple predicted maps based on observed information, and takes into consideration both the computed variances of predicted maps and estimated visible area to estimate the information gain of a given viewpoint. Experiments on the real-world KTH dataset showed on average 12.4% improvement than representative map-prediction based exploration and 25.4% improvement than nearest frontier approach.
Website: https://mapex-explorer.github.io/
@@ I Introduction
Exploration is an important problem in robotics, with applications ranging from indoor search and rescue operations [1] to planetary missions [2] and outdoor monitoring [3]. The primary objective of exploration is to plan paths in an unknown environment to maximize understanding of the environment within a limited budget. This objective may involve building an accurate map [4], identifying objects of interest [5], and finding traversable routes for robot teams [6]. The key challenge in exploration is choosing how to navigate to maximize information gain, which requires reasoning over unobserved space due to uncertain environment geometry, occlusions, sensor range limits, and sensor noise.
Figure 1: We present MapEx, a new framework that uses predicted maps for indoor exploration. Our key insight is to calculate probabilistic information gain that jointly reasons on coverage and uncertainty. We validate with a real-world floor plan dataset and show superior exploration performance compared to related map-prediction based exploration algorithms.
Many real-world environments possess inherent structure, predictability, and repeatability of environment geometry. For example, indoor environments such as offices or hospitals consist of repeated rooms and corridors, where observing partial views of the space can provide valuable information for predicting the overall layout. However, most existing approaches have difficulty leveraging this predictability for more efficient exploration. For instance, conventional frontier-based exploration [7, 8] do not directly reason over the predictable nature of the environment. Instead, they rely on heuristics such as “visit closest area first”.
To leverage such structural predictability, recent work has explored using deep learning techniques to predict unknown regions of maps, and use such prediction to improve exploration. For instance, IG-Hector [9] uses a predicted map to estimate how much area can be observed from a viewpoint. However, such viewpoint scoring heuristic is sensitive to the accuracy of the predicted map. UPEN [10] instead leverages such uncertain map prediction behavior, by explicitly predicting map uncertainty and then going towards these areas. However, this method does not consider sensor coverage or visibility, which do not reflect actual sensor acquisition. Therefore, our key insight is to jointly reason over what the robot can observe and its uncertainty to form probabilistic information gain from multiple predicted maps.
@@
  • Finally, we show the predicted maps produced by MapEx have improved global topological understanding and higher utility for downstream path planning.
@@ II-C Map Prediction for Exploration
Deep learning emerges as a promising approach for map prediction as the models and outputs are more expressive. Several works leverage deep learning-based map predictors with deep reinforcement learning-based planners for exploration [29, 30, 31]. However, these approaches are limited to short horizon decisions, such as 2.5m away from the robot [29], one step towards a direction [30, 31].
Our work builds on successful approaches that leverage deep-learning map prediction techniques and couple them with classical exploration planning techniques that allows more explicit planning over longer horizons, and no need for data to learn planning [9, 37, 28, 32, 10]. One method to using the prediction for exploration is to estimate the potential new information from a predicted sensor coverage at a given viewpoint and use it as an information gain metric [9, 37, 28, 32]. However, this information gain metric is susceptible to poor map predictions. Alternatively, UPEN [10] leverages such uncertainty in map prediction, to bias exploration towards high-uncertainty areas. UPEN estimates map uncertainty by predicting an ensemble of maps and using its variance as uncertainty. However, the information metric used is purely variance over the path and does not consider the potential sensor coverage, which is important to encapsulate actual information gain. To address the limitations of past work, MapEx jointly reasons over the predicted sensor coverage and the uncertainty of map prediction to form probabilistic information gain that is easily used by classical exploration frameworks [7], to generate long-horizon and informative exploration paths.
Figure 2: MapEx Framework: MapEx estimates probabilistic information gain from predicted maps to use in a frontier explorer. A robot accumulates observations into a top-down occupancy map, which is passed to an ensemble of map predictors to generate multiple map predictions. Information gain for each frontier is then estimated based on the mean and variance of the predicted maps. Finally, the frontier with the highest reward is set as the next goal.
@@ III Problem Statement
We address the problem of robot exploration in unknown indoor environments to gain a comprehensive understanding of the structure. Consider a robot exploring a 2D environment $\mathcal{E}\subset\mathcal{R}^{2}$. The robot’s state is represented as a 2D discrete state $\mathbf{x}=[x,y]$, and the robot’s action $a\in\mathcal{A}$ transitions its pose on a diagonally-connected grid at each timestep. The robot is equipped with a 2D LiDAR sensor with range $\lambda$. At time $t=0$, the robot starts with no prior knowledge of the environment, and at each timestep $t$, it receives 2D LiDAR measurements $o_{t}$ from evenly spaced rays spanning $360^{\circ}$ originating from the robot state $\mathbf{x_{t}}$, and takes an action $a_{t}$ according to the planner. We assume that the robot’s pose in a global coordinate frame is known and noise-free LiDAR range measurements are available. It continues to take actions and receive measurements until the time budget $T$.
@@ IV Approach
To efficiently explore structured indoor environments, we leverage predicted maps to identify areas with high probabilistic information gain. The MapEx pipeline, illustrated in Fig. 2, consists of four key steps:
@@ IV-A Global Map Prediction
Given the prevalence of large unknown areas during indoor exploration, we use LaMa network architecture [26], which has large receptive field and the ability to understand global context, making it suitable for inpainting maps with large unknown areas. We treat a top-down view occupancy grid as a 2D image (where 1 pixel = $0.1$m) and regard the unknown areas of the environment similarly to missing parts of an image. LaMa is used to predict what lies beyond the robot’s observation. To increase map prediction performance, we fine-tuned LaMa’s out-of-the-box Places2 weights using the KTH floor plan dataset[11]. Our training set consists of (observed map, ground truth map) data pairs collected through nearest-frontier exploration. Unlike prior works that predicts local maps [9, 10], we generate global predicted maps to enable more comprehensive global planning in exploration. Fig. 3 shows an example of an observed map, associated predicted map and ground-truth map.
@@ IV-B Uncertainty Estimation and Variance Map
While the robot can reasonably predict the map $P_{t}$ given the observed map $O_{t}$, there is inherent ambiguity in the unobserved areas. Therefore, the robot should estimate the uncertainty of the predictions. To achieve this, we maintain an ensemble of separate LaMa networks $\mathcal{G}_{i}$, and let each $\mathcal{G}_{i}$ generate a prediction:
@@ IV-C Calculating Probabilistic Information Gain
Given a viewpoint $c\in\mathcal{E}$, we generate $n_{l}$ hypothetical rays $R_{j}$ (where $j=1,2,\dots,n_{l}$) with a range of $\lambda$, emitted in a $360^{\circ}$ pattern originating from $c$. In a deterministic raycast approach, each ray $R_{j}$ stops when it hits an occupied cell, and if there are no obstructing cells along its path, $R_{j}$ reach the end at a distance of $\lambda$. In this work, we use a probabilistic raycast approach instead: first, we calculate the mean predicted map $\overline{P_{t}}$, which is the average of predicted maps $\{P_{i,t}\}$. Each ray $R_{j}$ starts with an initial accumulated occupancy value $\Delta=0$, and as $R_{j}$ passes through the pixels of $\overline{P_{t}}$, it adds the occupancy value of each pixel in the mean map $\overline{P_{t}}$ to $\Delta$. When $\Delta$ reaches a threshold $\epsilon$, the raycast terminates. Given ray end points, we perform a flood fill to generate a sensor coverage mask. We then mask out already observed areas from the sensor coverage mask to generate a visibility mask $\nu$, which probabilistically represents the unknown area estimated to be visible at location $c$. We then compute information gain I by summing the values of all pixels in the variance map $V_{t}$ that also fall within the visibility mask $\nu$. In other words,
@@ IV-D MapEx: Map Exploration with Prediction
Here, we explain how the MapEx algorithm is structured and uses the information gain values to augment a frontier-based explorer for efficient exploration. When selecting the next frontier, the robot first generates a predicted map $P_{t}$ based on the current observed occupancy grid $O_{t}$. Additionally, it generates an ensemble of predictions $\{P_{i,t}\}$, from which it computes the variance map $V_{t}$. The frontiers are extracted from the boundary between known and unknown cells in $O_{t}$, similar in [7]. We cluster frontiers using connected components, filter out small ones, and use the centroids of the remaining clusters as final frontiers. When calculating the score for these frontiers, the information gain is measured as described in Sec. IV-C. To discourage backtracking, we divide each frontier score by the Euclidean distance from the current pose of the robot to each frontier, so that if the scores are similar, the robot prefers to select a closer frontier. Once the waypoint is chosen, a local path is generated using an A* local planner, and the robot moves along the path. The pseudocode of the algorithm is provided in Alg. 1.
@@ V Experiments
We evaluate MapEx in a simulator using real-world floor plan datasets, comparing its performance against state-of-the-art baselines and conducting ablation studies.
@@ V-A Real-world Floor Plan Dataset
To benchmark real-world performance, we use the KTH floor plan dataset [11], which after removing repeated maps, contains 149 campus floor plan blueprints described as XML files of wall and door locations for each room. We adapt the processing code from [38] to obtain the occupancy map and manually correct inaccuracies in the processed map. We then downsample the map to $0.1$m per pixel. For training the map predictor and then testing the navigation, we split the floor plans into an 80:20 train-test split. We take care not to include floor plans from the same building into both the train and test set, to test the generalizability of our method.
@@ V-C Baselines
We also ablate the impact of different components of the information gain metric used by MapEx, to quantify the relative impact. For all ablation methods, we kept the frontier exploration framework the same, with only the frontier scoring method different. Specifically, we investigate
@@
  • No Visibility - No raycast performed. We sum the variance within $5$m of the viewpoint.
@@ V-D Metrics
We evaluate MapEx using a variety of metrics to comprehensively evaluate the information gained by the robot: coverage, predicted map quality, and topological understanding.
@@
Predicted Map Quality: The intersection over union (IoU) of the predicted 2D occupancy map and the ground truth occupancy map. Specifically, we use the IoU of the occupied class within the building footprint. A higher IoU indicates a more accurate predicted map.
Topological Understanding: We propose a new metric, Topological Understanding (TU), to directly evaluate how much the outcome of a robot’s exploration aids downstream path planning tasks. The TU metric is defined as follows: given the exploration result up to timestep $t$, represented by the predicted map $P_{t}$, we use it to plan an A* path from a start point to a goal. A path is considered successful if it reaches the goal. A path is considered a failure if it collides with an occupied cell in the ground truth map or fails to reach the goal. We measure success rates for 100 random goal locations per map across different methods.
@@ V-E1 Qualitative Analysis
Fig. 5 compares the exploration progress over time for a test map using different methods. UPEN [10] reasons on variance at the path and does not consider sensor coverage, likely leading it to explore areas with high variance near the center of the map. In contrast, at $t=500$, MapEx also explores the center, but quickly shifts to a more informative hallway, estimating high sensor coverage. IG-Hector [9], which focuses on sensor coverage, struggles early likely due to poor map predictions, causing inefficient exploration of tight spaces. In comparison, as seen in $t=200$MapEx searches a relatively open area, gathering information about surrounding rooms. Overall, MapEx explores more informative regions within the same timeframe, resulting in a higher-quality predicted map compared to the baselines. Fig. 6 provides additional qualitative examples of MapEx’s frontier scoring, showcasing the advantage of jointly reasoning on sensor coverage and uncertainty. Fig. 6a,b show top-scoring frontiers with high sensor coverage with medium uncertainty. In contrast, Fig. 6c presents a case where MapEx selects a frontier that covers smaller area but has high uncertainty, rather than one with higher coverage but less potential for learning, highlighting the importance of uncertainty-aware exploration.
@@ VII Acknowledgement
We thank Siddarth Narasimhan for sharing a more processed KTH dataset and providing valuable insights into baseline reproduction. We additionally thank Simon Stepputtis, Yaqi Xie, John Keller, David Fan and Mononito Goswami for helpful discussions.